# KACQ-DCNN: Uncertainty-Aware Interpretable Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network for Heart Disease Detection

### v1

@@ Abstract
Heart failure poses a critical global health challenge, significantly contributing to cardiovascular disease, which accounts for $\approx 17.8$ million deaths annually, underscoring the urgent need for innovative diagnostic strategies. Existing heart disease prediction models using classical machine learning techniques encounter significant challenges, particularly in managing complex, high-dimensional datasets with class imbalances, poor categorical feature representations, limited performance on small datasets, and a lack of uncertainty quantification in their predictions. Furthermore, these models often operate as ‘black boxes,’ making it difficult for healthcare professionals to interpret their outcomes, which hinders trust and clinical decision-making. Although recent advancements in quantum machine learning have shown great promise, current hybrid models that combine classical and quantum approaches have not yet fully leveraged the potential benefits of quantum computing. To address these gaps, we propose KACQ-DCNN, a novel classical–quantum hybrid dual-channel neural network that integrates Kolmogorov–Arnold Networks (KANs) in place of traditional multilayer perceptions and convolutional neural networks, enabling univariate learnable activation functions on edges. This approach enhances the ability of the network to approximate continuous functions with reduced complexity and improved generalizability. Our comprehensive evaluation demonstrates that the KACQ-DCNN 4-qubit 1-layered model outperforms 37 benchmark models, including 16 classical machine learning models, 12 quantum neural networks, six hybrid models, and three variants of the KACQ-DCNN architecture. The model achieved an accuracy of 92.03%, along with a macro-average precision, recall, and F1 score of 92.00%, representing significant improvements across all metrics. Moreover, the model achieved an ROC-AUC score of 94.77%, which was supported by two-tailed paired t-tests conducted against nine top-performing models, with a significance level ($\alpha$) of 5% and a Bonferroni correction applied ($\alpha_{\text{adjusted}}=0.0056$), using stratified 10-fold cross-validation. Next, we conducted ablation studies to demonstrate the synergistic effect of integrating classical–quantum components with KAN, highlighting their combined contribution to performance improvements. Furthermore, model-agnostic explainability techniques such as LIME and SHAP provide feature-level interpretability, enabling us to identify the most influential factors driving the model’s predictions and enhancing the transparency and trustworthiness of our KACQ-DCNN decision-making process. Uncertainty quantification, using conformal prediction techniques, has demonstrated robust performance and reliable probability estimates. Our findings suggest that the unique integration of the KACQ-DCNN model not only holds promise for more precise and timely interventions but also sets a foundation for future innovations that could revolutionize patient care and management in cardiovascular health.
@@
Classical
@@ Introduction
Traditional heart disease diagnostics rely on assessing risk factors such as age, sex, family history, lifestyle choices, and symptom analysis [4]. However, this manual approach often misses the hidden parameters and is computationally expensive [5, 2]. Although angiography remains the gold standard for diagnosing CAD [2, 4], its high cost and potential side effects limit its widespread use. Additionally, image-based diagnostic methods are impractical for large-scale screening because of their high cost. In response, researchers are increasingly focusing on developing noninvasive, cost-effective, and rapid automated diagnostic systems for early CAD detection via machine learning (ML) algorithms. Recent advances in artificial intelligence (AI) and ML have significantly improved diagnostic accuracy and accessibility by leveraging large datasets to uncover nonlinear and hierarchical patterns that traditional clinical methods often overlook. These technologies offer a more effective means of analyzing complex medical data and addressing the uncertainty and lack of interpretability that can lead to misdiagnoses or missed cases in conventional approaches.
@@
Current HF prediction models utilizing classical ML (CML) [1, 6, 7, 8, 9, 10, 11], deep learning (DL) [12, 4, 13, 14], hybrid [15, 16, 17], and quantum ML (QML) [2, 18, 19, 20, 21, 22, 23] techniques face several significant research gaps that limit their effectiveness, particularly when applied to short, imbalanced, and tabular datasets characterized by a mix of categorical and numerical features [24]. The integration of classical–quantum ML (CQML) techniques into heart disease prediction remains largely unexplored despite their potential to increase predictive performance through improved optimization and demonstrate computational capabilities even on short and imbalanced datasets [24, 25, 26]. Similarly, while the latest Kolmogorov–Arnold networks (KANs) [27, 28, 29] offer theoretical advantages over MLP-based models and CNNs—such as faster convergence using fewer parameters, learnable univariate functions on edges, optimization capability of activation functions, and smoother approximations with B-splines—they have yet to be fully implemented on practical datasets, limiting their real-world applications. Many ML-DL models operate as “black boxes,” making it challenging for healthcare professionals to understand the rationale behind predictions. This lack of transparency can hinder trust in these systems, especially when critical clinical decisions are at stake [30, 31]. Traditional models often fail to quantify the uncertainty in predictions, which is crucial for clinical decision-making. The integration of uncertainty-aware methods within classical–quantum frameworks could enhance the reliability of predictions; however, this has not been adequately explored. Heart disease datasets often suffer from class imbalance, where instances of heart disease are significantly fewer than those of non-disease cases. Traditional models may not adequately address this issue, leading to biased predictions that favor the majority class[32, 31]. Additionally, these models frequently struggle with overfitting, especially when trained on small datasets, leading to poor generalizability to unseen data[2, 33]. Furthermore, the handling of categorical features can be inadequate because traditional encoding methods may not effectively capture the underlying relationships. Traditional statistical methods often assume linear relationships between variables, which may not reflect the complex, nonlinear interactions present in cardiovascular health data[30, 33]. Moreover, the significant computational demands and extended training durations of DL models can hinder efforts to enhance the accuracy of heart sound classification [1].
@@
We present the “Kolmogorov–Arnold classical–quantum Dual-Channel Neural Network (KACQ-DCNN),” a pioneering hybrid model that combines the unique strengths of the KAN, CQML, explainable AI (XAI) methods such as SHAP and LIME, and uncertainty quantification through conformal prediction techniques. KACQ-DCNN addresses multiple challenges in ML, particularly in managing short, imbalanced, high-categorical, and high-dimensional datasets, offering a novel approach to solving real-world problems such as heart disease prediction. At the core of this innovation is the KAN-enhanced CQML model, which uses learnable activation functions on the edges, allowing it to approximate continuous functions with fewer parameters and reduced complexity. Its spline-based edge functions offer smoother and interpretable mappings, enabling faster convergence and improved generalization [28]. KAN excels in handling high-dimensional, nonlinearly separable data and provides the optimization facility of the activation function, unlike traditional MLPs and CNNs [29]. Furthermore, KACQ-DCNN leverages the exponential computational power of quantum techniques, allowing for the faster exploration of vast solution spaces and the ability to capture detailed correlations and patterns. The quantum components significantly boost scalability, making the model adept at handling both small and large datasets while retaining computational efficiency [24]. In addition, the incorporation of XAI tools such as SHAP and LIME enhances the model’s transparency, offering interpretable insights into predictions, which is especially valuable in fields such as healthcare. To ensure reliability, KACQ-DCNN integrates uncertainty quantification through conformal prediction, provides robust confidence intervals, and addresses ambiguous or high-risk classification scenarios.
@@
QML is an emerging field that combines quantum computing with traditional ML techniques, offering the potential to enhance computational power and improve diagnostic accuracy. Although still in its early stages, QML has shown potential for medical applications. Magallanes et al. [37] demonstrated the superiority of a hybrid classical–quantum model over CMLs in detecting stenosis from X-ray coronary angiography, significantly improving accuracy, recall, and F1 scores. Gupta et al. [38] compared QML techniques with DL models for diabetes prognosis using the PIMA Indian Diabetes dataset. Although the DL model achieved higher accuracy (95%), the QML model still performed well, showing promise in handling large datasets. Maheshwari et al. [39] explored the computational efficiency of quantum algorithms compared with classical algorithms in diabetes classification and reported that quantum approaches improved computational speed by up to 55 times. This suggests that although QML may not yet consistently outperform traditional models in terms of accuracy, its ability to handle vast datasets efficiently positions it as a valuable tool for medical prediction. Kumar et al. [40] compared traditional ML with QML algorithms and reported that the QML algorithm achieved an accuracy of 89%, with notable improvements in F1-score and recall. These findings suggest that although QML is still being developed, its advantages in handling large, complex datasets and improving computational efficiency are promising. QML has also gained traction; Kavitha and Kaulgud [41] presented a quantum K-means algorithm that attained 96.40% accuracy, marking significant progress in applying quantum computing to heart disease diagnostics. While current applications of QML are limited, the field is rapidly evolving, and future advancements may allow quantum models to surpass classical methods entirely.
@@
KANs represent the latest innovative approach to neural network architectures, offering enhanced accuracy with fewer parameters, dynamic adaptability, and architectural flexibility in medical diagnostics. Li et al. [42] developed the U-KAN model, a hybrid approach that combines KANs with U-Net for medical image segmentation tasks. This model reduces computational costs and improves segmentation performance, making it highly suitable for applications such as MRI analysis. In addition, KANs have been used for seizure detection based on EEG data. Contreras et al. [43] developed the KAN-EEG model, which demonstrated superior generalizability across datasets from different regions, reducing overfitting, and improving diagnostic accuracy. The flexibility and adaptability of KANs in medical diagnostics make them valuable, particularly for tasks requiring high interpretability and computational efficiency.
@@
ML has proven to be a valuable tool for heart disease prediction, with many models showing high accuracy and reliability. Although there are multiple databases related to heart disease, most studies have been conducted via the Cleveland database, and a few have used Statlog, demonstrating the need for more diverse and aggregated datasets. Conventional neural networks typically demand many parameters and extensive computational resources, posing a significant challenge in environments with limited resources. Traditional ML algorithms such as RF and SVM have demonstrated robustness in clinical settings, while hybrid and DL models offer enhanced accuracy for complex datasets. However, previous research focused only on a single type of dataset at a time, which might contain potential bias, lack of generalizability, or overfitting because of being short, highly categorical, and imbalanced. The emerging field of QML holds promise for improving computational efficiency and diagnostic performance, although it has yet to outperform classical methods consistently. KANs, as substitutes for MLP, also offer a novel approach to improving the accuracy with less computational complexity and flexibility of neural network architectures in medical diagnostics. Integrating these advanced techniques will likely lead to more precise, efficient, and interpretable models for heart disease prediction.
@@ Methods
MERA [44, 45] is a sophisticated tensor network structure designed to efficiently capture and represent the entanglement properties of quantum many-body systems, particularly those with scale-invariant or critical behavior. Unlike more superficial tensor network structures such as MPS or TTN, MERA incorporates disentangling operations known as isometries, which systematically reduce correlations at different scales. This hierarchical approach allows MERA to represent quantum states with long-range entanglement, thus making it a powerful tool for studying critical phenomena in quantum systems. We employed the MERA structure to model the complex correlations between the dataset features. Our implementation uses a variational quantum classifier (VQC) framework, where the MERA layers act as entangling blocks of the quantum circuit. Specifically, the quantum circuit comprises a series of $CNOT$ gates and $RY$ rotations controlled by trainable parameters. In quantum computing, the CNOT gate is a 2-qubit gate that flips the state of the target qubit if the control qubit is in the $|1\rangle$ state and can be represented as a $4\times 4$ unitary matrix:
@@
MPS [46] represents a powerful and versatile tensor network approach for modeling quantum many-body systems. In the MPS framework, the quantum state of a one-dimensional system is expressed as the product of matrices, each associated with a site in the chain. This representation is beneficial for capturing local correlations and is foundational for developing algorithms such as density matrix renormalization group (DMRG).
@@
Integrating the BiLSTMKANnet and QDenseKANnet models is designed as a dual-channel architecture that leverages both classical DL and quantum-enhanced computations. The architecture begins by processing the input data separately through two channels: classical BiLSTMKANnet and hybrid quantum-classical QDenseKANnet. Both models accept input tensors of shape $(\text{None},12,1)$, meaning samples with 12 features, each of them represented as a single value. and 1 can be interpreted as the feature dimension, commonly used in neural networks to handle input with multiple features in a consistent shape. However, the data undergo distinct transformations within each network, owing to their strengths.
@@
For stratified 10-fold cross-validation, all models went through 10 rounds of experiments, each comprising a split, training, and testing phase. In each round, the dataset was divided into 10 equal folds, with 9 folds used for training and 1 fold for testing. This process was repeated 10 times to ensure that each fold was used for testing exactly once. The reported ‘$\pm$’ values in Table 3 and 5 represent the mean and standard deviation of the performance metrics across these 10 rounds, providing a robust measure of both the average performance and the variability for each model.
@@
In contrast to the existing literature, mentioned in Table 4, where models were trained on individual heart disease datasets, our approach is novel in that we merged multiple datasets, including the Statlog and UCI dataset, to show generalizability and robustness of KACQ-DCNN. This is the first attempt in the literature to train a model using such a merged heart disease dataset, offering a more holistic view across different datasets. It’s important to highlight that the Cleveland dataset is a small subset of the UCI dataset (Cleveland, Hungarian, Switzerland, and Long Beach, Virginia), on which several studies trained models individually. Our method allows for broader model generalization, potentially capturing patterns across datasets, which could improve robustness in real-world applications. Despite this complexity, our proposed KACQ-DCNN achieved 92% across all evaluation metrics, indicating strong performance even when dealing with the diversity of combined datasets. In the literature, dataset-specific models were used to achieve high accuracy using only one dataset. For example, Mir et al.[16] got 99.48% accuracy via RF + Adaboost on the Heart Statlog dataset. However, KACQ-DCNN also achieved superior state-of-the-art performance on Cleveland and Statlog datasets, both sharing an equal maF1 of 96% and the accuracy of 96.23% and 96.30%, respectively.
@@
Over 16 epochs, KACQ-DCNN initially showed consistent improvements in terms of loss and accuracy. The first epoch resulted in a loss of 0.7087, with a validation accuracy of 73.55%. By the third epoch, validation accuracy reached 88.77%, and by the seventh epoch, it peaked at 92.03%. The accuracy subsequently fluctuated without significant improvements. To enhance convergence, the learning rate was reduced in epoch 12 from $1\times 10^{-3}$ to $5\times 10^{-4}$. However, validation accuracy could not surpass the 92.03% peak, and training was halted at epoch 16 owing to the early stopping mechanism. As shown in Figure LABEL:fig:loss, the training loss steadily decreased across all epochs, indicating effective learning. However, the validation loss plateaued around epoch 7, with minor fluctuations suggesting that the model was near its generalization capacity. Despite the learning rate reduction at epoch 12, validation loss showed no substantial gains, confirming the onset of overfitting. As depicted by Figure LABEL:fig:cm, KACQ-DCNN correctly predicts 40.22% of normal cases and 51.81% of HF cases, indicating good performance in detecting HFs. However, 4.35% of normal cases were misclassified as HF, whereas 3.62% of HF cases were incorrectly labeled as normal.
@@
Figures LABEL:fig:roc_curve and LABEL:fig:pr_curve illustrate the performance of KACQ-DCNN via the ROC and precision-recall curves, which are critical metrics for evaluating binary classification tasks. The ROC curve demonstrates a high TP rate against the FP rate, indicating that the model effectively distinguishes between the positive and negative classes. An AUC score of 99.73% on the training set and 94.77% on the test set suggested excellent classification ability, with minimal risk of overfitting. Similarly, the precision-recall curve reveals the model’s ability to maintain high precision and recall rates, achieving values of 99.78% in training and 94.89% in testing. This finding indicates that the model is not only accurate but also robust across both datasets, demonstrating its effectiveness in predicting the target classes.
@@
KACQ-DCNN 1-Layered 92.00%$\pm$0.1992.00%$\pm$0.1092.00%$\pm$0.0792.03%$\pm$0.1994.77%$\pm$0.4383.85%$\pm$0.8183.84%$\pm$0.43497.64 10.49

### v3

@@ Abstract
Heart failure remains a critical global health issue, contributing significantly to cardiovascular disease and accounting for $\approx 17.8$ millions of annual deaths. The need for innovative diagnostic strategies is pressing, as classical machine learning models face challenges such as handling complex, high-dimensional data, class imbalances, poor categorical feature representations, limited performance on small datasets, and the absence of uncertainty quantification. Moreover, the interpretability of these models is often hindered by their ‘black box’ nature, complicating clinical trust and decision-making. While quantum machine learning shows potential, existing hybrid models have yet to fully capitalize on quantum advantages. To address these gaps, we propose Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network (KACQ-DCNN), a novel hybrid dual-channel neural network that integrates Kolmogorov–Arnold Networks (KANs) in place of traditional multilayer perceptions, enabling univariate learnable activation functions on edges. As an early adopter of KAN components, we observed that the approach significantly improved the ability of the network to approximate continuous functions with reduced complexity and improved generalizability. Our comprehensive evaluation demonstrates that the KACQ-DCNN 4-qubit 1-layered model outperforms 37 benchmark models, including 16 classical machine learning models, 12 quantum neural networks, six hybrid models, and three variants of the KACQ-DCNN architecture. It achieved an accuracy of 92.03%, along with a macro-average precision, recall, and F1 score of 92.00%, representing significant improvements across all metrics. Moreover, KACQ-DCNN achieved a ROC-AUC score of 94.77%, supported by two-tailed paired t-tests against nine top-performing models, with a significance level ($\alpha$) of 5% and a Bonferroni correction applied ($\alpha_{\text{adjusted}}=0.0056$), using stratified 10-fold cross-validation. Ablation studies confirmed the synergistic effect of integrating classical-quantum components with KAN, improving accuracy by $\approx 2\%$ compared to MLP variants. Furthermore, model-agnostic explainability techniques like LIME and SHAP enhanced feature-level interpretability, and conformal prediction techniques provided robust uncertainty quantification. Our findings indicate that KACQ-DCNN advances prediction accuracy for timely interventions and provides a foundation for more interpretable and trustworthy cardiovascular diagnostics.
@@
Kolmogorov–Arnold Networks (KAN), Classical-Quantum Neural Network, Conformal Prediction, Uncertainty Quantification, Explainable AI, Heart Disease Prediction
@@ Introduction
Traditional heart disease diagnostics rely on assessing risk factors such as age, sex, family history, lifestyle choices, and symptom analysis [4]. However, this manual approach often fails to capture the hidden parameters and is computationally expensive [5, 2]. Although angiography remains the gold standard for diagnosing CAD [2, 4], its high cost and potential side effects limit its widespread use. Additionally, image-based diagnostic methods are impractical for large-scale screening because of their high cost. In response, researchers are increasingly focusing on developing noninvasive, cost-effective, and rapid automated diagnostic systems for early CAD detection via machine learning (ML) algorithms. Recent advances in artificial intelligence (AI) and ML have significantly improved diagnostic accuracy and accessibility by leveraging large datasets to uncover nonlinear and hierarchical patterns that traditional clinical methods often overlook. These technologies offer a more effective means of analyzing complex medical data and addressing the uncertainty and lack of interpretability that can lead to misdiagnoses or missed cases in conventional approaches.
@@
Current HF prediction models utilizing classical ML (CML) [1, 6, 7, 8, 9, 10, 11], deep learning (DL) [12, 4, 13, 14], hybrid [15, 16, 17], and quantum ML (QML) [2, 18, 19, 20, 21, 22, 23, 24] techniques face several significant research gaps that limit their effectiveness, particularly when applied to short, imbalanced, and tabular datasets characterized by a mix of categorical and numerical features [25]. The integration of classical–quantum ML (CQML) techniques into heart disease prediction remains largely unexplored despite their potential to increase predictive performance through improved optimization and demonstrate computational capabilities even on short and imbalanced datasets [25, 26, 27]. Similarly, while the latest Kolmogorov–Arnold networks (KANs) [28, 29, 30, 31] offer theoretical advantages over MLP-based models and CNNs—such as faster convergence using fewer parameters, learnable univariate functions on edges, optimization capability of activation functions, and smoother approximations with B-splines—they have yet to be fully implemented on practical datasets, limiting their real-world applications such as heart failure detection and investigating its interdisciplinary combinations with fields like QML.
Many ML-DL models operate as “black boxes,” making it challenging for healthcare professionals to understand the rationale behind predictions. This lack of transparency can hinder trust in these systems, especially when critical clinical decisions are at stake [32, 33]. Traditional models often fail to quantify the uncertainty in predictions, which is crucial for clinical decision-making. The integration of uncertainty-aware methods within classical–quantum frameworks could enhance the reliability of predictions; however, this has not been adequately explored. Heart disease datasets often suffer from class imbalance, where instances of heart disease are significantly fewer than those of non-disease cases. Traditional models may not adequately address this issue, leading to biased predictions that favor the majority class[34, 33]. Additionally, these models frequently struggle with overfitting, especially when trained on small datasets, leading to poor generalizability to unseen data[2, 35]. Furthermore, handling categorical features can be inadequate because traditional encoding methods may not effectively capture the underlying relationships. Traditional statistical methods often assume linear relationships between variables, which may not reflect the complex, nonlinear interactions present in cardiovascular health data[32, 35]. Moreover, the significant computational demands and extended training durations of DL models can hinder efforts to enhance the accuracy of heart sound classification [1].
@@
We present the “Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network (KACQ-DCNN),” a hybrid model that combines the unique strengths of the KAN, CQML, explainable AI (XAI) methods such as SHAP and LIME, and uncertainty quantification through conformal prediction techniques. KACQ-DCNN addresses multiple challenges in ML by replacing MLP layers with KAN layers, particularly in managing short, imbalanced, high-categorical, and high-dimensional datasets, offering a novel approach to solving real-world problems such as heart disease prediction. At the core of this innovation is the KAN-enhanced CQML model, which uses learnable activation functions on the edges, allowing it to approximate continuous functions with fewer parameters and reduced complexity. Its spline-based edge functions offer smoother and interpretable mappings, enabling faster convergence and improved generalization [29]. KAN excels in handling high-dimensional, nonlinearly separable data and provides the optimization facility of the activation function, unlike traditional MLPs and CNNs [30, 31]. Furthermore, KACQ-DCNN leverages the exponential computational power of quantum techniques, allowing for the faster exploration of vast solution spaces and the ability to capture detailed correlations and patterns. The quantum components significantly boost scalability, making the model adept at handling small and large datasets while retaining computational efficiency [25]. We selected BiLSTMKANnet as channel-1 and QDenseKANnet as channel-2 after thorough benchmarking, as both models demonstrated superior performance compared to all other evaluated options. Our choice of a dual-channel neural network (DCNN) was driven by its ability to process diverse types of input data in parallel, leveraging distinct feature extraction capabilities for each channel. This approach allows us to combine the individual predictions from each channel to produce final probabilities, effectively enhancing overall performance without compromising efficiency. This architecture enabled us to effectively capture both local and global contexts, which previously worked well with sentiment analysis tasks [36], where understanding nuanced interactions is crucial. Furthermore, it minimizes the risk of overfitting and offers robustness against noise and variability in real-world applications, allowing the model to adapt to dynamic data streams while maintaining accuracy [37, 38]. In addition, incorporating XAI tools such as SHAP and LIME enhances the model’s transparency, offering interpretable insights into predictions, which is especially valuable in fields such as healthcare. To ensure reliability, KACQ-DCNN integrates uncertainty quantification through conformal prediction, provides robust confidence intervals, and addresses ambiguous or high-risk classification scenarios.
@@
QML is an emerging field that combines quantum computing with traditional ML techniques, offering the potential to enhance computational power and improve diagnostic accuracy. Although still in its early stages, QML has shown potential for medical applications. Magallanes et al. [42] demonstrated the superiority of a hybrid classical–quantum model over CMLs in detecting stenosis from X-ray coronary angiography, significantly improving accuracy, recall, and F1 scores. Gupta et al. [43] compared QML techniques with DL models for diabetes prognosis using the PIMA Indian Diabetes dataset. Although the DL model achieved higher accuracy (95%), the QML model still performed well, showing promise in handling large datasets. Maheshwari et al. [44] explored the computational efficiency of quantum algorithms compared with classical algorithms in diabetes classification and reported that quantum approaches improved computational speed by up to 55 times. This suggests that although QML may not yet consistently outperform traditional models in terms of accuracy, its ability to handle vast datasets efficiently positions it as a valuable tool for medical prediction. Kumar et al. [23] compared traditional ML with QML algorithms and reported that the QML algorithm achieved an accuracy of 89% in heart failure detection, with notable improvements in F1-score and recall. These findings suggest that although QML is still being developed, its advantages in handling large, complex datasets and improving computational efficiency are promising. QML has also gained traction; Kavitha and Kaulgud [24] presented a quantum K-means algorithm that attained 96.40% accuracy, marking significant progress in applying quantum computing to heart disease diagnostics. While current applications of QML are limited, the field is rapidly evolving, and future advancements may allow quantum models to surpass classical methods entirely.
@@
KANs represent the latest innovative approach to neural network architectures, offering enhanced accuracy with fewer parameters, dynamic adaptability, and architectural flexibility in medical diagnostics. Li et al. [45] developed the U-KAN model, a hybrid approach that combines KANs with U-Net for medical image segmentation tasks. This model reduces computational costs and improves segmentation performance, making it highly suitable for applications such as MRI analysis. In addition, KANs have been used for seizure detection based on EEG data. Contreras et al. [46] developed the KAN-EEG model, which demonstrated superior generalizability across datasets from different regions, reducing overfitting and improving diagnostic accuracy. The flexibility and adaptability of KANs in medical diagnostics make them valuable, particularly for tasks requiring high interpretability and computational efficiency.
@@
ML has proven to be a valuable tool for heart disease prediction, with many models showing high accuracy and reliability. Although there are multiple databases related to heart disease, most studies have been conducted via the Cleveland database, and a few have used Statlog, demonstrating the need for more diverse and aggregated datasets. Conventional neural networks typically demand a higher number of parameters and extensive computational resources, posing a significant challenge in environments with limited resources. Traditional ML algorithms such as RF and SVM have demonstrated robustness in clinical settings, while hybrid and DL models offer enhanced accuracy for complex datasets. However, previous research focused only on a single type of dataset at a time, which might contain potential bias, lack of generalizability, or overfitting because of being short, highly categorical, and imbalanced. While previous studies explored QML and CQML in heart disease prediction [2, 24, 23], showing their promise for improving computational efficiency and diagnostic performance, they have yet to outperform classical methods consistently. KANs, as substitutes for MLP, offer a novel approach to improving the accuracy with less computational complexity and flexibility of neural network architectures in medical diagnostics, which remains greatly unexplored. Integrating KANs with CQML techniques will likely lead to more precise, efficient, and interpretable models for heart disease prediction.
@@ Methods
MERA is a sophisticated tensor network structure designed to efficiently capture and represent the entanglement properties of quantum many-body systems, particularly those with scale-invariant or critical behavior. Unlike more superficial tensor network structures such as MPS or TTN, MERA incorporates disentangling operations known as isometries, which systematically reduce correlations at different scales. This hierarchical approach allows MERA to represent quantum states with long-range entanglement, thus making it a powerful tool for studying critical phenomena in quantum systems. We employed the MERA structure to model the complex correlations between the dataset features. Our implementation uses a variational quantum classifier (VQC) framework, where the MERA layers act as entangling blocks of the quantum circuit. Specifically, the quantum circuit comprises a series of $CNOT$ gates and $RY$ rotations controlled by trainable parameters. In quantum computing, the CNOT gate is a 2-qubit gate that flips the state of the target qubit if the control qubit is in the $|1\rangle$ state and can be represented as a $4\times 4$ unitary matrix:
@@
MPS represents a powerful and versatile tensor network approach for modeling quantum many-body systems. In the MPS framework, the quantum state of a one-dimensional system is expressed as the product of matrices, each associated with a site in the chain. This representation is beneficial for capturing local correlations and is foundational for developing algorithms such as density matrix renormalization group (DMRG).
@@
Integrating the BiLSTMKANnet and QDenseKANnet models is designed as a dual-channel architecture that leverages both classical DL and quantum-enhanced computations. The architecture begins by processing the input data separately through two channels: classical BiLSTMKANnet and hybrid quantum-classical QDenseKANnet. Both models accept input tensors of shape $(\text{None},12,1)$, meaning samples with 12 features, each of them represented as a single value. and 1 can be interpreted as the feature dimension, commonly used in neural networks to handle input with multiple features in a consistent shape. However, due to their strengths, the data undergo distinct transformations within each network.
@@
For stratified 10-fold cross-validation, all models went through 10 rounds of experiments, each comprising a split, training, and testing phase. The dataset was divided into 10 equal folds in each round, with 9 folds used for training and 1 fold for testing. This process was repeated 10 times to ensure that each fold was used for testing exactly once. The reported ‘$\pm$’ values in Table 3 and 5 represent the mean and standard deviation of the performance metrics across these 10 rounds, providing a robust measure of both the average performance and the variability for each model.
@@
In contrast to the existing literature, mentioned in Table 4, where models were trained on individual heart disease datasets, our approach is novel in that we merged multiple datasets, including the Statlog and UCI dataset, to show generalizability and robustness of KACQ-DCNN. This is the first attempt in the literature to train a model using such a merged heart disease dataset, offering a more holistic view across different datasets. It’s important to highlight that the Cleveland dataset is a small subset of the UCI dataset (Cleveland, Hungarian, Switzerland, and Long Beach, Virginia), on which several studies trained models individually. Our KAN-enhanced method allows for broader model generalization, potentially capturing patterns across datasets, which could improve robustness in real-world applications. Despite this complexity, our proposed KACQ-DCNN achieved 92% across all evaluation metrics, indicating strong performance even when dealing with the diversity of combined datasets. In the literature, dataset-specific models were used to achieve high accuracy using only one dataset. For example, Mir et al. [16] got 99.48% accuracy via RF + Adaboost on the Heart Statlog dataset. However, KACQ-DCNN also achieved superior state-of-the-art performance on Cleveland and Statlog datasets, both sharing an equal maF1 of 96% and the accuracy of 96.23% and 96.30%, respectively.
@@
Over 16 epochs, KACQ-DCNN initially showed consistent improvements in terms of loss and accuracy. The first epoch resulted in a loss of 0.7087, with a validation accuracy of 73.55%. By the third epoch, validation accuracy reached 88.77%, and by the seventh epoch, it peaked at 92.03%. The accuracy subsequently fluctuated without significant improvements. The learning rate was reduced in epoch 12 to enhance convergence from $1\times 10^{-3}$ to $5\times 10^{-4}$. However, validation accuracy could not surpass the 92.03% peak, and training was halted at epoch 16 owing to the early stopping mechanism. As shown in Figure LABEL:fig:loss, the training loss steadily decreased across all epochs, indicating effective learning. However, the validation loss plateaued around epoch 7, with minor fluctuations suggesting that the model was near its generalization capacity. Despite the learning rate reduction at epoch 12, validation loss showed no substantial gains, confirming the onset of overfitting. As depicted by Figure LABEL:fig:cm, KACQ-DCNN correctly predicts 40.22% of normal cases and 51.81% of HF cases, indicating good performance in detecting HFs. However, 4.35% of normal cases were misclassified as HF, whereas 3.62% of HF cases were incorrectly labeled as normal.
@@
Figures LABEL:fig:roc_curve and LABEL:fig:pr_curve illustrate the performance of KACQ-DCNN via the ROC and precision-recall curves, which are critical metrics for evaluating binary classification tasks. The ROC curve demonstrates a high TP rate against the FP rate, indicating that the model effectively distinguishes between the positive and negative classes. An AUC score of 99.73% on the training set and 94.77% on the test set suggested excellent classification ability, with minimal risk of overfitting. Similarly, the precision-recall curve reveals the model’s ability to maintain high precision and recall rates, achieving values of 99.78% in training and 94.89% in testing. This finding indicates that the model is accurate and robust across both datasets, demonstrating its effectiveness in predicting the target classes.
@@
KAN version of KACQ-DCNN (1-Layered) 92.00%$\pm$0.1992.00%$\pm$0.1092.00%$\pm$0.0792.03%$\pm$0.1994.77%$\pm$0.4383.85%$\pm$0.8183.84%$\pm$0.43497.64 10.49